# Automatic Mapping of AutomationML Files to Ontologies for Graph Queries and Validation

### v1

@@ Abstract
AutomationML has seen widespread adoption as an open data exchange format in the automation domain. It is an open and vendor neutral standard based on the extensible markup language XML. However, AutomationML extends XML with additional semantics, that limit the applicability of common XML-tools for applications like querying or data validation. This article provides practitioners with 1) an up-to-date ontology of the concepts in the AutomationML-standard, as well as 2) a declarative mapping to automatically transform any AutomationML model into RDF triples. Together, these artifacts allow practitioners an easy integration of AutomationML information into industrial knowledge graphs. A study on examples from the automation domain concludes that transforming AutomationML to OWL opens up new powerful ways for querying and validation that are impossible without transformation.
@@ 1 Introduction
AutomationML is designed to represent complex relationships within engineering projects. It is used to model and exchange data related to the design, implementation, and maintenance of automation systems. The AutomationML data model takes the form of a deeply interconnected graph. However, for easier exchange with other engineering tools, AutomationML uses XML as its underlying format for serialization, which follows a tree structure. In the serialization, references between elements of the XML-tree are realized through reference paths and Identifiers.
In practice, this representation mechanism offers multiple drawbacks. First, common tools to query and validate XML-Files (like XSD or xQuery) were designed to operate on the hierarchical structure of the data format. They therefore have very rudimentary support for operations involving the additional connections defined by AutomationML. Second, the data lacks a formal representation which would allow knowledge inference and further analytical applications[3]. Third, the XML-Schema of AutomationML does not allow for native alignment to other data sources, i.e. it is limited to the vocabulary that AutomationML has to offer. However, this alignment is a pillar of the digital twin[2], making it especially necessary in the engineering domain [4].
In terms of formally representing engineering knowledge, technologies like the Resource Description Framework (RDF) and the Web Ontology Language (OWL), are widely used, as they provide a standardized framework for expressing relationships and semantics [5]. These technologies have been applied to create meta models of cyber-physical systems [6][7], which allows for the integration of heterogeneous data and resolving semantic heterogeneity [4]. RDF and OWL further allow the use of technologies from the Semantic Web Stack like SPARQL and SHACL, which further enable efficient querying and validation of the created knowledge graphs.
@@
However, each of these proposed AutomationML ontologies exhibits significant shortcomings in their scope, design methodology, or compliance with current AutomationML specifications. Additionally, there is no freely available tool or mapping definition to automatically transform an AutomationML file into an RDF graph, further limiting widespread adoption.
Thus, this article aims to close these gaps by providing practitioners with two artifacts to be reused in future works:
1.
An up to date AutomationML Ontology, with full coverage of the standards vocabulary.
2.
Using these artifacts, it is possible to automatically translate an AutomationML file into an ontology. This translation process enables new use cases for querying and validating AutomationML files, which are demonstrated on an example system.
The remainder of the article is structured as follows: Section 2 gives more information on the AutomationML format, how it relates to graph theory and on the tools from the Semantic Web Stack. Section 3 describes previous attempts at an AutomationML ontology, while Section 4 discusses the necessary changes and additions to the AutomationML ontology. Section 5 describes the design decisions of mapping AutomationML files to RDF. The mapping’s correctness and flexibility are then validated on multiple AutomationML files in Section 6. Here, the AutomationML files are used for various query and data validation scenarios. The use cases show that transforming AutomationML to RDF allows for more complex graph queries and validation rules in comparison to typical XML-based approaches.
@@ 2.1 AutomationML
AutomationML (Automation Markup Language) is an open, XML-based standard for the exchange of engineering data. It builds on established and standardized data exchange formats like Computer Aided Engineering Exchange (CAEX) for plant topology, COLLADA for geometry and kinematics, and PLCopenXML for control logic. It is commonly used in both the scientific community and industry[1, 12] and has subsequently been standardized in IEC 62714[13].
@@
The InstanceHierarchy is a hierarchy of InternalElements and contains the data of a specific project.
The SystemUnitClassLibrary provides SystemUnitClasses, which describe types of physical plant objects, e.g. a specific type of robot or valve.
The RoleClassLibrary provides RoleClasses, which define abstract roles that a physical or logical plant objects may fill. While SystemUnitClasses usually describe some manufacturer specific systems, RoleClasses model the abstract functions of components.
The InterfaceClassLibrary provides InterfaceClasses which define a type of interface (e.g. a flange or a signal interface). These interfaces are used to describe relationships between InternalElements (IE) and SystemUnitClasses (SUC) through InternalLinks.
The AttributeTypeLibrary defines a predefined data structure of reusable AttributeTypes and can be used to model attributes with values, units and descriptions. It allows for the definition of proprietary or standardized vocabularies of attributes, and creates a basis for the semantic interpretability of attributes.
The content of an AutomationML file is structured in the shape of an XML-hierarchy, in which each element has a specific path. Additionally, AutomationML makes intensive use of references and links between different elements on different branches on this tree. These references between elements of the XML-tree are realized through reference paths. Using this mechanism, classes and roles can be linked to instances of this class. It further allows for the definition of interfaces, which function as connection points between different AutomationML elements.
@@ 2.2 AutomationML and Graph Theory
XML-based files (like AutomationML) can be considered as a specific subtype of graphs. Here, the XML-Elements are considered as the nodes, whereas the hierarchical organization of elements forms the edges. If viewed in this way, XML files naturally take the shape of a Tree (s. left side of Figure 2). Trees are a specific type of acyclic, connected graph:
@@
XML’s hierarchical nature and the unique path from the root to each element ensure that the data can be traversed, queried, and manipulated as a tree, making it an ideal format for structured data representation. XML-based tools for querying and validation heavily rely on this hierarchical parent / child structure.
However, as stated in Section 2.1, AutomationML also allows additional references between nodes besides their hierarchical position in the XML file. Once the ReferencePaths and InternalLinks are considered, the AutomationML model no longer satisfies the property of having a unique path between every pair of nodes. An example of such a cyclical AutomationML graph can be seen on the right side of Figure 2. It is therefore no longer a Tree as defined by Definition 2
@@
XML tools for queries (like xPath and xQuery) as well as for validation (like XSD and Schematron) were designed to handle hierarchical data in the shape of Trees (s. Equation 2). They therefore only offer limited support for connections that are not part of the XML tree structure. For AutomationML, this includes any connection that involves referencing another element via either reference path or ID, i.e. any class assignment or InternalLink between elements.
@@ 2.3 Semantic Web Technologies
Semantic Web Technologies are a framework of technologies designed to enable machines to meaningfully interpret and process data. The Resource Description Framework (RDF), and its Schema Extension (RDFS) and the Web Ontology Language (OWL) are related technologies used in the context of the Semantic Web Technologies. They are often used together to represent and reason about knowledge in a machine-readable format.
@@
The Web Ontology Language OWL, on the other hand, is a language for defining ontologies. An ontology is a formal description of concepts, relationships, and properties within a certain domain[16]. OWL allows for more expressive and complex modeling of relationships between entities than RDF: While RDF provides a foundation for expressing relationships, OWL enhances the expressiveness and formalism of these relationships, enabling the creation of ontologies for more advanced knowledge representations. In combination with a reasoner, OWL enables the inference of additional knowledge.
@@ 3 Requirements and State of the Art
In the past, numerous attempts have been made to create a formalized AutomationML ontology. The following Section first outlines the requirements that a usable AutomationML ontology needs to fulfill. Afterwards it gives an overview of related works and previous AutomationML ontologies, before comparing previous ontologies to the requirements.
@@ 3.1 Requirements
The resulting knowledge graph should be both highly expressive as well as easy to use for people that are already familiar with AutomationML’s terminology. This results in a number of requirements, that influence later design decisions. For instance, in order to make the ontology easy to understand and use for AutomationML-practitioners, new terminology should be avoided, resulting in Requirement R1. Since queries as well as validation use cases make use of the entire width of information that is stored in an AutomationML file, the ontology should likewise cover the entire vocabulary that CAEX has to offer (R2). However, the CAEX vocabulary within AutomationML is subject to contextual interpretation based on hierarchical positioning and inter-element relationships. For example the RefBaseSystemUnitClass-Property can be used to define the SystemUnitClass of an InternalElement. Used in a different context however, it can also describe a Mirror/Master relationship between two InternalElements (s. Sec. 5.7). Recognizing these contextual semantics and making them explicitly available are important for both accurate modeling as well as usability of the resulting knowledge graph (R3).
AutomationML is an actively maintained and developed data exchange format. The ontology should therefore consider the most recent version of the AutomationML specification (R4).
Most applications of Semantic Web Technologies make heavy use of OWL terminology concepts like class/property hierarchies and property characteristics. The ontology should therefore enrich and align AutomationML concepts with the terminology used in OWL (R5).
Manually translating an AutomationML file to OWL involves extensive effort and can introduce ambiguities in the mapping process. In order to allow practitioners to actually use the ontology beyond the scope of prototypical implementations, a reusable method to automatically translate AutomationML files to ontologies is needed (R6).
@@
3.2 Related Works: Mapping and Applications
Sabou et al. [21] present the AutomationML-Analyzer, which is a tool that showcases the functionalities made possible by the combination of AutomationML files with semantic web technologies. The authors use the aforementioned AutomationML Ontology presented in [22]. The mapping is performed using a two-step process, in which the AutomationML-files are first converted to Ecore, a domain model of the Eclipse Modeling Framework, and then converted from Ecore to OWL. This conversion procedure is identified as less than ideal by the authors themselves, as they propose a direct mapping as a potential future work. With regards to querying AutomationML-files, Wimmer et al. [23] define an AutomationML-specific query language. While it expands upon the capabilities of other XML-based Query languages, it similarly uses the Ecore-Mapping.
The formal representation of the knowledge in an AutomationML file has been used as the basis for numerous applications over the past years. For example, Abele et al. [8] transformed CAEX-models of production plants to OWL and subsequently used the resulting formal representation to perform a design validation using semantic web reasoning.
Grangel-Gonzales et al. [10] used knowledge from AutomationML files to semantically describe cyber-physical systems from different design perspectives. Gonçalves et al. [11] use the engineering knowledge contained in an AutomationML file with an ontology for sensor fusion to deal with the occurrence of data divergence.
@@ 3.3 Prior AutomationML-Ontologies
In the past, multiple attempts at an AutomationML ontology have been published. Abele et al.[8] cover the most important aspects of the CAEX vocabulary, and even some basic context-dependent concepts like MirrorObjects. However, neither of those are complete as lots of additional information about the entities is not covered. An automated mapping approach using Apache Jena is mentioned, but unavailable.
Grangel-Gonzales et al. [9] focus on the Attribute structure of AutomationML and are therefore lacking coverage of the remaining CAEX vocabulary. AutomationML files are mapped automatically using a XSLT-based tool. Due to the limited coverage of the model, the mapping method is not viable for large scale projects.
Kovalenko et al. [22] focus on the description of the CAEX vocabulary, but without covering the Attributes. An automated mapping approach using Apache Jena is mentioned, but unavailable.
@@
Since the release of all considered ontologies, the AutomationML standard underwent continuous development and has since been released as Version 3.0 in 2022[14]. All ontologies were released before this date and can therefore not fulfill Requirement R4. Because of this, they lack concepts for the semantic AttributeLibraries introduced in [14].
@@
R5: Enrich Model with OWL concepts | ||||
R6: Automated Mapping Approach |
As Table 2 shows, none of the considered Ontologies fulfill all requirements. However, the ontology presented by Kovalenko et al. [3] achieves good coverage of the basic CAEX concepts. Since the new AutomationML version[14] introduces significant changes to the AutomationML vocabulary as well as the XML serialization, the decision was made to develop a new ontology. Since many the basic CAEX concepts did not change, and [3] achieved good coverage in that area, their ontology was used as a starting point.

### v2

# utomationML]xml frame=lines, bgcolor=gray!10, fontsize=, tabsize=2 PARQL]sparql frame=lines, bgcolor=gray!10, fontsize=, tabsize=2 Path]xpath frame=lines, bgcolor=gray!10, fontsize=, tabsize=2 Query]xquery frame=lines, bgcolor=gray!10, fontsize=, tabsize=2 Automatic Mapping of AutomationML Files to Ontologies for Graph Queries and Validation

@@ Abstract
AutomationML has seen widespread adoption as an open data exchange format in the automation domain. It is an open and vendor neutral standard based on the extensible markup language XML. However, AutomationML extends XML with additional semantics that limit the applicability of common XML-tools for applications like querying or data validation. This article demonstrates how the transformation of AutomationML into OWL enables new use cases in querying with SPARQL and validation with SHACL. To support this, it provides practitioners with (1) an up-to-date ontology of the concepts defined in the AutomationML standard and (2) a declarative mapping to automatically transform any AutomationML model into RDF triples. A study on examples from the automation domain concludes that transforming AutomationML to OWL opens up new powerful ways for querying and validation that would have been impossible without this transformation.
@@ 1 Introduction
AutomationML is designed to represent complex relationships within engineering projects. It is used to model and exchange data related to the design, implementation, and maintenance of automation systems. The AutomationML data model takes the form of a deeply interconnected graph. However, for easier exchange with other engineering tools, AutomationML uses XML as its underlying format for serialization, which follows a tree structure. In the serialization, references between AutomationML elements of the XML-tree are realized through reference paths and identifiers.
In practice, this representation mechanism offers multiple drawbacks. First, commonly used tools to query and validate XML-Files (like XSD or xQuery) were designed to operate on the hierarchical structure of the data format. They therefore have very rudimentary support for operations involving the additional connections defined by AutomationML. Second, the AutomationML data lacks a formal representation which would allow knowledge inference and further analytical applications[3]. Third, the XML-Schema of AutomationML does not allow for native alignment with other data sources, i.e. it is limited to the vocabulary that AutomationML has to offer. However, this alignment is a pillar of the digital twin[2], making it especially necessary in the engineering domain [4].
In terms of formally representing engineering knowledge, technologies like the Resource Description Framework (RDF) and the Web Ontology Language (OWL) are widely used, as they provide a standardized framework for expressing relationships and semantics [5]. These technologies have been applied to create meta models of cyber-physical systems [6][7], which allows for the integration of heterogeneous data and resolving semantic heterogeneity [4]. RDF and OWL further allow the use of technologies from the Semantic Web Stack like SPARQL and SHACL, which further enable efficient querying and validation of the created knowledge graphs.
@@
However, none of the existing approaches have systematically examined the extent to which such a transformation enables new applications. While prior work has primarily emphasized data integration, the potential benefits of representing AutomationML as an ontology—such as the ability to formulate expressive semantic queries or to apply logical validation techniques—remain largely unexplored.
To address this gap, the paper compares the querying capabilities of XML-based tools with those of SPARQL, and examines how validation in XML-based environments contrasts with the possibilities offered by SHACL. In doing so, it introduces an up-to-date AutomationML ontology that covers the full vocabulary of the standard, presents a declarative mapping that enables the straightforward transformation of AutomationML files into RDF triples and provides a software implementation of this mapping to support its practical application. Together, these contributions provide practitioners with both the technical foundations and practical means to exploit ontology-based querying and validation in AutomationML beyond the limits of existing XML-based approaches.
The remainder of the article is structured as follows: Section 2 gives more information on the AutomationML format, how it relates to graph theory and the tools from the Semantic Web Stack. Section 3 describes previous attempts to create an AutomationML ontology and a mapping, while Section 4 presents the newly proposed ontology. Section 5 formalizes the mapping rules to transform AutomationML files to OWL, which are exemplified in Section 6. Section 7 evaluates the approach by comparing querying with XML-based tools and SPARQL, as well as validation with XML-based mechanisms and SHACL. The use cases show that transforming AutomationML to RDF allows for more complex graph queries and validation rules in comparison to typical XML-based approaches.
@@ 2.1 AutomationML
AutomationML (Automation Markup Language) is an open, XML-based standard for the exchange of engineering data. It builds on established and standardized data exchange formats like Computer Aided Engineering Exchange (CAEX) for plant topology[12], COLLADA for geometry and kinematics [13], and PLCopenXML for control logic[14]. It is commonly used in both the scientific community and industry[1, 15] and has subsequently been standardized in IEC 62714[16].
@@
The InstanceHierarchy is a hierarchy of InternalElements (IE) and contains the data of a specific project.
The SystemUnitClassLibrary provides SystemUnitClasses (SUC), which describe types of physical plant objects, e.g. a specific type of robot or valve.
The RoleClassLibrary provides RoleClasses (RC), which define abstract roles that a physical or logical plant objects may fill. While SystemUnitClasses usually describe some manufacturer specific systems, RoleClasses model the abstract functions of components.
The InterfaceClassLibrary provides InterfaceClasses (IC) which define a type of interface (e.g. a flange or a signal interface). These interfaces are used to describe physical or logical connections between InternalElements (IE) and SystemUnitClasses (SUC) through InternalLinks.
The AttributeTypeLibrary defines a predefined data structure of reusable AttributeTypes (AT) and can be used to model attributes with values, units and descriptions. It allows for the definition of proprietary or standardized vocabularies of attributes, and creates a basis for the semantic interpretability of attributes.
The content of an AutomationML file is structured in the shape of an XML-hierarchy. However, the semantics of this hierarchy can vary depending on the context within the AutomationML structure. For example, in the instance hierarchy, parent-child relationships represent physical or logical containment. In contrast, within class libraries such as RoleClassLibraries, AttributeTypeLibraries and InterfaceClassLibraries, the hierarchy is primarily used for structuring and organizing the data. Notably, it does not imply inheritance or specialization by default, but rather provides a logical grouping of related elements for better readability and maintainability.
Beyond the hierarchical organization, AutomationML supports multiple referencing mechanisms to relate elements across the file. Most references – such as those used for class inheritance, interface instantiation, and role or class assignments – are based on a reference path system. These reference paths point to elements within the structured hierarchy using relative or absolute identifiers, allowing connections to be expressed in a way that reflects the modular and nested organization of the XML structure.
In contrast, InternalLinks are used to explicitly define connections between Interface elements by referencing their unique IDs. Since Interfaces typically represent defined interaction points—such as mechanical, electrical, or software interfaces—InternalLinks provide a way to model physical or logical connections between them. This mechanism enables the representation of cross-hierarchy associations that are not captured by structural containment alone.
Listing 2 shows examples of hierarchy-, path- and ID-based references.
<InstanceHierarchy Name="MyInstanceHierarchy"> <InternalElement Name="MixingPlant1"> <InternalElement Name="Pipe1"> <ExternalInterface Name="Output" ID="6eea7a40-43fd-4aee-a1d3"> </ExternalInterface> </InternalElement> <InternalElement Name="Tank1"> <ExternalInterface Name="Input" ID="ce45d7b3-169d-4be8-9746"> </ExternalInterface> <Attribute Name="Length" RefAttributeType="MyAtLib/Dimensions/Length"> </Attribute> <InternalLink Name="Pipe1_to_Tank1" RefPartnerSideA="6eea7a40-43fd-4aee-a1d3" RefPartnerSideB="ce45d7b3-169d-4be8-9746"/> </InternalElement> </InternalElement></InstanceHierarchy>
AutomationML adopts the CAEX data model and terminology, but some terms acquire context-dependent meanings within AutomationML. Certain AutomationML specific concepts, such as Groups or Facets, are technically modeled as InternalElement objects with a predefined RoleRequirement, thereby specializing standard CAEX constructs. In other cases, semantics shift depending on the connection type—for example, a RefBaseSystemUnitClass link may either indicate an association between an InternalElement and a SystemUnitClass, or represent a master/mirror relationship between two InternalElements.
@@ 2.2 AutomationML and Graph Theory
XML-based files can be considered as a specific subtype of graphs. Here, the XML-Elements are considered as the nodes, whereas the hierarchical organization of elements forms the edges. If viewed in this way, XML files naturally take the shape of a Tree (s. solid black lines of Figure 3). Trees are a specific type of acyclic, connected graph:
@@
XML’s hierarchical nature ensure that the data can be traversed, queried, and manipulated as a tree, making it an ideal format for structured data representation. XML-based tools for querying and validation heavily rely on this hierarchical parent / child structure.
However, as stated in Section 2.1, AutomationML also allows additional references between nodes besides their hierarchical position in the XML file. Once the ReferencePaths and InternalLinks are considered, the AutomationML model no longer satisfies the property of having a unique path between every pair of nodes. An example of such a AutomationML graph can be seen in Figure 3. It is therefore no longer a Tree as defined by Definition 2
@@
XML tools for queries (like xPath and xQuery) as well as for validation (like XSD and Schematron) were designed to handle hierarchical data in the shape of Trees (s. Equation 2). They therefore only offer limited support for connections that are not part of the XML tree structure (i.e. all dashed lines in Figure 3).
@@ 2.3 Semantic Web Technologies
Semantic Web Technologies are a framework of technologies designed to enable machines to meaningfully interpret and process data. The Resource Description Framework (RDF), its Schema Extension (RDFS) and the Web Ontology Language (OWL) are related technologies used in the context of the Semantic Web Technologies. They are often used together to represent and reason about knowledge in a machine-readable format.
@@
The Web Ontology Language OWL, on the other hand, is a language for defining ontologies. An ontology is a formal description of concepts, relationships, and properties within a certain domain[19]. While RDF provides a foundation for expressing relationships, OWL enhances the expressiveness and formalism of these relationships, enabling the creation of more advanced knowledge representations.
@@ 3 Requirements and State of the Art
In the past, numerous attempts have been made to create an AutomationML ontology. The following Section first outlines the requirements that a usable AutomationML ontology needs to fulfill. Afterwards it gives an overview of related works and previous AutomationML ontologies, before comparing previous ontologies to the requirements.
@@ 3.1 Requirements
The resulting knowledge graph should be both highly expressive as well as easy to use for people that are already familiar with AutomationML’s terminology. This results in a number of requirements, that influence later design decisions. In order to make the ontology easy to understand and use for AutomationML-practitioners, the terminology of the ontology should closely follow the established AutomationML terminology, resulting in Requirement R1.. Since queries as well as validation use cases make use of the entire width of information that is stored in an AutomationML file, the ontology should likewise cover the entire vocabulary that AutomationML has to offer (R2.). However, as stated in Section 2.1, the vocabulary within AutomationML is subject to contextual interpretation based on hierarchical positioning and inter-element relationships. Recognizing these contextual semantics and making them explicitly available are important for both accurate modeling as well as usability of the resulting knowledge graph (R3.).
AutomationML is an actively maintained and developed data exchange format. The ontology should therefore consider the most recent version of the AutomationML specification (R4.).
Most applications of Semantic Web Technologies make heavy use of OWL terminology concepts like class/property hierarchies and property characteristics. Besides following the AutomationML terminology, the ontology should therefore supplement the AutomationML concepts with the logically interpretable vocabulary used by OWL (R5.).
The terminological definitions in an AutomationML ontology usually consist of two parts: First, the ’core’ AutomationML terminology, that defines the AutomationML metamodel (e.g. what a RoleClass is). This is needed to formally describe any AutomationML file. The second part is the individual classification hierarchy (e.g. which RoleClasses there are). These hierarchies change between projects, and can therefore not be modeled overarchingly for all AutomationML files. Since some translation effort for the creation and population of the ontology is therefore project specific, modeling ’just’ the core AutomationML terminology in an ontology is not sufficient. Creating a full AutomationML ontology for a project thus necessitates a reusable mapping procedure (R6.).
@@ 3.2 Prior AutomationML-Ontologies
In the past, multiple attempts to create an AutomationML ontology have been published.
Runde et al.[23] were the first to show the potential of combining CAEX with semantic web technologies. Their approach focuses mostly on representing CAEX using OWL concepts. However, the approach is lacking with regard to the representation of attributes, and deviates from the terminology that CAEX/AutomationML establish.
Abele et al.[8] cover the most important aspects of the CAEX vocabulary, and even some basic context-dependent concepts like MirrorObjects. However, neither of those are complete as lots of additional information about the entities is not covered.
Grangel-Gonzales et al. [9] focus on the Attribute structure of AutomationML and are therefore lacking coverage of the remaining CAEX vocabulary. AutomationML files are mapped automatically using a XSLT-based tool.
Kovalenko et al. [24] focus on the description of the AutomationML vocabulary, but the authors do not completely cover the full AutomationML schema specification, particularly for data properties.
@@
Since the release of all considered ontologies, the AutomationML standard underwent continuous development and has since been released as Version 3.0 in 2022[17]. All ontologies were released before this date and can therefore not fulfill Requirement R4.. Because of this, they lack concepts for the semantic AttributeLibraries introduced in [17].
@@
R5: Enrich Model with OWL concepts |
As Table 2 shows, none of the considered Ontologies fulfill all requirements. However, the ontology presented by Kovalenko et al. [3] achieves good coverage of the basic CAEX concepts. Since the new AutomationML version[17] introduces significant changes to the AutomationML vocabulary as well as the XML serialization, a new ontology needed to be developed. Since many of the basic CAEX concepts did not change, and [3] achieved good coverage in that area, their ontology was used as a starting point.
Section 4 describes the content of this newly developed ontology.
3.3 Prior Mapping Approaches
Multiple publications approached the translation of AutomationML to OWL, using different target ontologies as well as translation technologies.
Out of the ontologies mentioned in Section 3.2, Runde et al.[23], Kovalenko et al.[24] and Grangel-Gonzales et al.[9]mention an automatic translation. Runde et al.[23] propose a semi-automated mapping approach, but do not mention the implementation technology. Kovalenko et al. [24] propose a transformation from AutomationML into OWL using the software framework Apache Jena, which is unavailable. Finally, Grangel-Gonzales et al.[9] propose translating AutomationML to RDF using Krextor, an XSLT-based framework for converting XML to RDF. However, Krextor has not been maintained.
Sabou et al. [25] present the AutomationML-Analyzer, which is a tool that showcases the functionalities made possible by the combination of AutomationML-files with semantic web technologies. The authors use the aforementioned AutomationML Ontology presented in [24]. The mapping is performed using a two-step process, in which the AutomationML-files are first converted to Ecore, a domain model of the Eclipse Modeling Framework, and then converted from Ecore to OWL. This conversion procedure is identified as less than ideal by the authors, as they propose a direct mapping as a potential future work.
Approach | Technology | Target Ontology | Available |
Runde et al.[23] | - | [23] | |
Kovalenko et al.[24] | Apache Jena | [24] | |
Grangel-Gonzales et al.[9] | Krextor/XSLT | [9] | |
Sabou et al.[25] | Ecore/Java | [24] | |
Hua et al.[26] | Custom/Java | NA |
The preexisting mapping definitions target differ in their underlying technology as well as their target ontologies. As stated in Section 3.2, these ontologies to not fully meet the outlined requirements, and will therefore be replaced by a new AutomationML ontology. Since the publication of these mapping approaches, RML has emerged as the de facto standard declarative mapping language to automatically convert structured data into RDF triples, making it a natural choice for the implementation of a new mapping. Sections 5 and 6 describe the mapping rules for the transformation as well as a detailed example.